# Communication Cost Reduction for Subgraph Counting under Local Differential Privacy via Hash Functions

### v1

@@ IV-D Counting Step
Recall that $d_{i}$ is the number of connections with nodes of smaller indexes of user $v_{i}$, which will be called low degree from here after. Let $p_{\text{p},i}:=\mathbb{P}\left(a_{j,h_{j}(k)}=1|a_{j,k}=1\right)$, i.e.
$p_{\text{p},i}=\frac{s-1}{s}\cdot\frac{d_{i}-1}{ms-1}+\frac{1}{s},$
and let $p^{\prime}_{\text{p},i}:=\mathbb{P}\left(a^{\prime}_{j,h_{j}(k)}=1|a_{j,k}=1\right)$, i.e.
@@
Also, let $p_{\text{a},i}:=\mathbb{P}\left(a_{j,h_{j}(k)}=1|a_{j,k}=0\right)$, i.e.
$p_{\text{a},i}=\frac{s-1}{s}\cdot\frac{d_{i}}{ms-1},$
and let $p^{\prime}_{\text{a},i}:=\mathbb{P}\left(a^{\prime}_{j,h_{j}(k)}=1|a_{j,k}=0\right)$, i.e.
@@
Define
$\omega_{i}$ ::=:= $\frac{1}{p_{\text{p},i}^{\prime}-p_{\text{a},i}^{\prime}}$
$=$ $1/\left(\frac{e^{\varepsilon^{\prime}}-1}{e^{\varepsilon^{\prime}}+1}(p_{\text{p},i}-p_{\text{a},i})\right)$
$=$ $1/\left(\frac{e^{\varepsilon^{\prime}}-1}{e^{\varepsilon^{\prime}}+1}\left(\frac{1}{s}-\frac{1}{ms-1}\right)\right)$
$=$ $\frac{e^{\varepsilon^{\prime}}+1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms^{2}-s}{ms-s-1}.$
Also, define
$\sigma_{i}$ ::=:= $\frac{p_{\text{a},i}^{\prime}}{p_{\text{p},i}^{\prime}-p_{\text{a},i}^{\prime}}$
$=$ $\left(\frac{e^{\varepsilon^{\prime}}-1}{e^{\varepsilon^{\prime}}+1}\cdot p_{\text{a},i}+\frac{1}{1+e^{\varepsilon^{\prime}}}\right)\cdot\frac{e^{\varepsilon^{\prime}}+1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms^{2}-s}{ms-s-1}$
$=$ $p_{\text{a},i}\cdot\frac{ms^{2}-s}{ms-s-1}+\frac{1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms^{2}-s}{ms-s-1}$
$=$ $\frac{s-1}{ms-s-1}d_{i}+\frac{1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms^{2}-s}{ms-s-1}.$
Recall that $\tilde{d}_{i}$ is the low degree estimation of user $v_{i}$ published in the first step of our framework. Define
$\tilde{\sigma}_{i}:=\frac{s-1}{ms-s-1}\tilde{d}_{i}+\frac{1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms^{2}-s}{ms-s-1}.$
We have that $\mathbb{E}[\tilde{\sigma}_{i}]=\sigma_{i}$. For all $i,t$, denote
$\tilde{a}_{i,c_{t}}:=\omega_{i}a_{i,c_{t}}^{\prime}-\tilde{\sigma}_{i}.$
In the previous work, subgraphs are counted based on the value of $a_{j,k}^{\prime}\in\{0,1\}$ obtained from the randomized mechanism. In GRR, we replace $a_{j,k}^{\prime}$ with $\tilde{a}_{j,c_{t}}$ when $t=h_{j}(k)$. Let $W_{i}:=\{(j,k):j<k<i\text{ and }\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E\}$. The estimated number of triangles (denoted by $|\Delta_{i}|$ was $\sum\limits_{(j,k)\in W_{i}}a^{\prime}_{j,k}$ in the previous work. In this work, the estimated number of triangle corresponding to $v_{i}$, denoted by $|\tilde{\Delta}_{i}|$, is $\sum\limits_{(j,k)\in W_{i}}\tilde{a}_{j,h_{j}(k)}$.
Similar to the previous method, each user calculate the number of subgraphs corresponding to their nodes. They then report the number using the Laplacian mechanism or the smooth sensitivity mechanism with the privacy budget $\varepsilon_{2}$. The estimated number of subgraphs is then equal to the sums of those reported from users.
V Theoretical Analysis of Our Mechanism
V-A Privacy
The privacy of our mechanism, discussed in the next theorem, is a direct result of the theorem on amplification by sub-sampling [34].
Theorem 2.
[…]
As the randomized response step has a privacy budget of $\varepsilon^{\prime}=\ln(1+s(e^{\epsilon_{1}}-1))$, the release of $a_{i,c_{1}},\dots,a_{i,c_{m}}$ in the first two steps has a privacy budget of $\ln(1+(e^{\varepsilon^{\prime}}-1)/s)=\ln(1+(1+s(e^{\varepsilon_{1}}-1)-1)/s)=\varepsilon_{1}$.
Since users do not release any information during the grouping step, the privacy budget of our four-step mechanism is then equal to $\varepsilon_{0}+\varepsilon_{1}+\varepsilon_{2}$ by the composition theorem.
V-B Expected Value of $\tilde{a}_{i,c_{t}}$
In the subsequent theorem, we discuss the expected value of $\tilde{a}_{i,c_{t}}$. The theorem reveals that when an estimate from the counting phase is a linear composition of $\tilde{a}_{i,c_{t}}$, it remains unbiased. The triangle count estimation in Section 4.3 is an example of these estimations.
@@ V-B Expected Value of $\tilde{a}_{i,c_{t}}$
When $a_{j,k}=0$,
$\mathbb{E}\left[\tilde{a}_{j,h_{j}(k)}|a_{j,k}=0\right]$
$=$ $(\omega_{j}-\mathbb{E}[\tilde{\sigma}_{j}])\cdot p^{\prime}_{\text{a},j}-\mathbb{E}[\tilde{\sigma}_{j}]\cdot(1-p^{\prime}_{\text{a},j})$
$=$ $\omega_{j}p_{\text{a},j}^{\prime}-\sigma_{j}=0.$
When $a_{j,k}=1$, we have that:
$\mathbb{E}\left[\tilde{a}_{j,h_{j}(k)}|a_{j,k}=1\right]$
$=$ $(\omega_{j}-\mathbb{E}[\tilde{\sigma}_{j}])\cdot p^{\prime}_{\text{p},j}-\mathbb{E}[\tilde{\sigma}_{j}]\cdot(1-p^{\prime}_{\text{p},j})$
$=$ $\omega_{j}p^{\prime}_{\text{p},j}-\sigma_{j}=1.$
V-C Variance of $\tilde{a}_{i,c_{t}}$
@@
When $a_{j,k}=0$, we have that $\mathbb{P}\left(a^{\prime}_{j,h_{j}(k)}=1\right)=p^{\prime}_{\text{a},j}=\frac{\sigma_{j}}{\omega_{j}}.$ When $a_{j,k}=1$, we have that $\mathbb{P}\left(a^{\prime}_{j,h_{j}(k)}=1\right)=p^{\prime}_{\text{p},j}=\frac{1+\sigma_{j}}{\omega_{j}}.$
@@
By $\varepsilon^{\prime}=\ln(1+s(e^{\varepsilon_{1}}-1))$, we have
$e^{\varepsilon^{\prime}}-1$ $=$ $s(e^{\varepsilon_{1}}-1)$
$\frac{1}{e^{\varepsilon^{\prime}}-1}$ $=$ $\frac{1}{s(e^{\varepsilon_{1}}-1)}.$
Hence,
$\omega_{i}$ $=$ $\frac{2+s(e^{\varepsilon_{1}}-1)}{s(e^{\varepsilon_{1}}-1)}\cdot\frac{ms^{2}-s}{ms-s-1}$
$=$ $O\left(\frac{s}{e^{\varepsilon_{1}}-1}\right).$
By the assumption that $d_{i}<m$, we have
$\sigma_{i}$$\leq$ $\frac{1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms^{2}-s}{ms-s-1}+1$
$=$ $\frac{1}{s(e^{\varepsilon_{1}}-1)}\cdot\frac{ms^{2}-s}{ms-s-1}+1$
$\leq$ $1+\frac{m}{m-1}\cdot\frac{1}{e^{\varepsilon_{1}}-1}=O\left(\frac{1}{e^{\varepsilon_{1}}-1}\right).$
@@
We first consider the variance of the variable $a^{\prime}_{j,h_{j}(k)}$. Because the variable is binary, by Lemma 1, we obtain that
$\text{Var}\left(a^{\prime}_{j,h_{j}(k)}\right)$ $=$ $\mathbb{P}\left(a^{\prime}_{j,h_{j}(k)}=1\right)\cdot\left(1-\mathbb{P}\left(a^{\prime}_{j,h_{j}(k)}=1\right)\right)$
$=$ $\left(\frac{a_{j,k}+\sigma_{j}}{\omega_{j}}\right)\cdot\left(1-\frac{a_{j,k}+\sigma_{j}}{\omega_{j}}\right)$
$\leq$ $\frac{1+\sigma_{j}}{\omega_{j}}.$
Then, by Lemma 2,
$\text{Var}\left(\tilde{a}_{j,h_{j}(k)}\right)$ $=$ $\text{Var}\left(\omega_{j}a^{\prime}_{j,h_{j}(k)}+\tilde{\sigma}_{i}\right)$
$=$ $\omega_{j}^{2}\text{Var}\left(a^{\prime}_{j,h_{j}(k)}\right)+\text{Var}\left(\tilde{\sigma}_{i}\right)$
$\leq$ $\omega_{j}(1+\sigma_{j})+\left(\frac{s-1}{ms-s-1}\right)^{2}\frac{2}{\varepsilon_{0}^{2}}$
$=$ $O\left(\frac{s}{(e^{\varepsilon_{1}}-1)^{2}}+\frac{s^{2}}{n^{2}\varepsilon_{0}^{2}}\right).$
In the following subsection, we will elaborate that $\tilde{a}_{j,h_{j}(k)}$ can cut the communication cost by a factor of $s^{2}$ in contrast to using the direct randomized response results $a^{\prime}_{j,k}$. Given that the variance of $a^{\prime}_{j,k}$ is known to be $\Theta\left(\frac{1}{(e^{\epsilon_{1}}-1)^{2}}\right)$, the previous theorem suggests that, while we manage to decrease the communication cost by a factor of $s^{2}$, this is at the expense of amplifying the variance by a factor of $s$. In the upcoming section, we will demonstrate that, for a set communication cost, this mechanism yields a considerably reduced variance.
In situations where the goal is to sum up specific estimators, like in subgraph counting, a covariance emerges between these estimators. The size of this covariance is addressed in the subsequent theorem:
@@
When $j\neq j^{\prime}$, the two estimators are sampled from different sets of variables. They are thus independent.
@@
When $j=j^{\prime}$, the two estimators are identical when $k$ and $k^{\prime}$ are in the same bin, i.e. $h_{j}(k)=h_{j}(k^{\prime})$. This event, that we will denote $C^{(j)}_{k,k^{\prime}}$, happens with a probability of $\frac{s-1}{ms-1}$. As a result, for $k,k^{\prime}<j$:
$\text{Cov}\left(\tilde{a}_{j,h_{j}(k)},\tilde{a}_{j,h_{j}(k^{\prime})}\right)$
=\ = $\omega_{j}^{2}\cdot\text{Cov}\left(a^{\prime}_{j,h_{j}(k)},a^{\prime}_{j,h_{j}(k^{\prime})}\right)$
@@
=\ = $\omega_{j}^{2}\cdot\left(\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\right]-\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}\right]\mathbb{E}\left[a^{\prime}_{j,h_{j}(k^{\prime})}\right]\right).$
Using those notations, we will consider the values of $\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\mid C^{(j)}_{k,k^{\prime}}\right]$ and $\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\mid\overline{C^{(j)}_{k,k^{\prime}}}\right]$.
Recall from the previous section that $\mathbb{P}\left(a_{j,h_{j}(k)}=1\right)=\frac{s-1}{s}\cdot\frac{d_{j}}{ms-1}+a_{j,k}\left[\frac{1}{s}-\frac{1}{ms-1}\right]$. We have that:
$\frac{ms-1}{ms-2}\mathbb{P}\left(a_{j,h_{j}(k)}=1\right)$
$=$ $\frac{s-1}{s}\cdot\frac{d_{i}}{ms-2}+a_{j,k}\left[\frac{ms-1}{s(ms-2)}-\frac{1}{ms-2}\right]$
$\geq$ $\frac{s-1}{s}\cdot\frac{d_{i}-a_{j,k}}{ms-2}+\frac{a_{j,k}}{s}.$
For $\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\mid\overline{C^{(j)}_{k,k^{\prime}}}\right]$, let us denote $\mathbb{P}\left(a_{j,h_{j}(k)}=1\mid\overline{C^{(j)}_{k,k^{\prime}}}\right)$ as $p_{k}$. We obtain that:
$p_{k}$ $=$ $\frac{a_{j,k}}{s}+\frac{s-1}{s}\cdot\frac{d_{j}-a_{j,k}-a_{j,k^{\prime}}}{ms-2}$
$\leq$ $\frac{ms-1}{ms-2}\cdot\mathbb{P}\left(a_{j,h_{j}(k)}=1\right).$
Using those notations, we have:
$\mathbb{P}\left(a^{\prime}_{j,h_{j}(k)}=1\text{ and }a^{\prime}_{j,h_{j}(k^{\prime})}=1\mid\overline{C^{(j)}_{k,k^{\prime}}}\right)$
=\ = $\left(p_{k}\frac{e^{\varepsilon^{\prime}}-1}{e^{\varepsilon^{\prime}}+1}+\frac{1-p_{k}}{e^{\varepsilon^{\prime}}+1}\right)\left(p_{k^{\prime}}\frac{e^{\varepsilon^{\prime}}-1}{e^{\varepsilon^{\prime}}+1}+\frac{1-p_{k^{\prime}}}{e^{\varepsilon^{\prime}}+1}\right)$
[…]
In consequence, as long as $s$ is larger than 3, which will always be the case in practice, we have that:
@@
$\mathbb{P}\left(\overline{C^{(j)}_{k,k^{\prime}}}\right)\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\mid\overline{C^{(j)}_{k,k^{\prime}}}\right]\\ -\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}\right]\mathbb{E}\left[a^{\prime}_{j,h_{j}(k^{\prime})}\right]\leq 0$
This simplifies the computation as we now have that:
@@
$\text{Cov}\left(\tilde{a}_{j,h_{j}(k)},\tilde{a}_{j,h_{j}(k^{\prime})}\right)\\ \leq\omega_{j}^{2}\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\mid C^{(j)}_{k,k^{\prime}}\right]\mathbb{P}\left(C^{(j)}_{k,k^{\prime}}\right)$
Now, let us denote $\mathbb{P}\left(a_{j,h_{j}(k)}=1\mid C^{(j)}_{k,k^{\prime}}\right)$ as $p^{\prime}_{k}$. We have that:
$p_{k}^{\prime}$ $=$ $\frac{a_{j,k}+a_{j,k^{\prime}}}{s}+\frac{s-2}{s}\cdot\frac{d_{j}-a_{j,k}-a_{j,k^{\prime}}}{ms-2}$
$\leq$ $\mathbb{P}\left(a_{j,h_{j}(k)}=1\right)+\mathbb{P}\left(a_{j,h_{j}(k^{\prime})}=1\right).$
Using the same argument as in the previous paragraph, we obtain that:
@@
By Lemma 1 and the fact that $\mathbb{P}\left(C_{k,k^{\prime}}^{(j)}\right)=\frac{s-1}{ms-1}$, we get:
$\text{Cov}\left(\tilde{a}_{j,h_{j}(k)},\tilde{a}_{j,h_{j}(k^{\prime})}\right)$$\leq$ $2\frac{s-1}{ms-1}\omega_{j}(1+\sigma_{j})$
$=$ $O\left(\frac{s^{2}}{n(e^{\varepsilon_{1}}-1)^{2}}\right).$
It can be noticed that, when $s\ll n$, the covariance in the previous theorem is much smaller than the variance calculated in Theorem 4. We thus believe that the variance dominates almost all the loss we have found in the experiment.
V-D Communication Cost

### v2

@@ 5.2 Expected Value of $\tilde{a}_{i,c_{t}}$
Recall that $d_{i}$ denotes the number of connections between user $v_{i}$ and nodes with smaller indices, which we refer to as the low degree of $v_{i}$.
Define
$p_{\text{p},i}:=\Pr[a_{j,h_{j}(k)}=1\mid a_{j,k}=1],$
which evaluates to
$p_{\text{p},i}=\frac{s-1}{s}\cdot\frac{d_{i}-1}{ms-1}+\frac{1}{s}.$
Let
$p^{\prime}_{\text{p},i}:=\Pr[a^{\prime}_{j,h_{j}(k)}=1\mid a_{j,k}=1],$
which is given by the randomized response mechanism as
@@
Similarly, define
$p_{\text{a},i}:=\Pr[a_{j,h_{j}(k)}=1\mid a_{j,k}=0]=\frac{s-1}{s}\cdot\frac{d_{i}}{ms-1},$
and let
$p^{\prime}_{\text{a},i}:=\Pr[a^{\prime}_{j,h_{j}(k)}=1\mid a_{j,k}=0],$
which satisfies
@@
In the next lemma, we show a property of $\omega_{i}$ and $\tilde{\sigma}_{i}$ calculated in Algorithm 3. From now, denote $\mathbb{E}[\tilde{\sigma}_{i}]$ by $\sigma_{i}$.
Lemma 3.
For all $i$, $\omega_{i}=\frac{1}{p_{\text{p},i}^{\prime}-p_{\text{a},i}^{\prime}}$ and $\sigma_{i}=\frac{p_{\text{a},i}^{\prime}}{p_{\text{p},i}^{\prime}-p_{\text{a},i}^{\prime}}$.
Proof.
We have
$\frac{1}{p_{\text{p},i}^{\prime}-p_{\text{a},i}^{\prime}}=\frac{1}{\left(\frac{e^{\varepsilon^{\prime}}-1}{e^{\varepsilon^{\prime}}+1}\right)(p_{\text{p},i}-p_{\text{a},i})}=\frac{1}{\left(\frac{e^{\varepsilon^{\prime}}-1}{e^{\varepsilon^{\prime}}+1}\right)\left(\frac{1}{s}-\frac{s-1}{s}\cdot\frac{1}{ms-1}\right)}.$
Simplifying the expression gives
$\frac{1}{p_{\text{p},i}^{\prime}-p_{\text{a},i}^{\prime}}=\frac{e^{\varepsilon^{\prime}}+1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms-1}{m-1}=:\omega_{i}.$
On the other hand, we obtain
$\frac{p_{\text{a},i}^{\prime}}{p_{\text{p},i}^{\prime}-p_{\text{a},i}^{\prime}}=\left(\frac{s-1}{ms-s}\cdot d_{i}\right)+\left(\frac{1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms-1}{m-1}\right)=\left(\frac{s-1}{ms-s}\cdot\mathbb{E}[\tilde{d}_{i}]\right)+\left(\frac{1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms-1}{m-1}\right)=:\sigma_{i}.$
We are now ready the proof the main theorem of this subsection.
@@
When $a_{j,k}=0$, the expected value of $\tilde{a}_{j,h_{j}(k)}$ is given by
$\mathbb{E}\left[\tilde{a}_{j,h_{j}(k)}\mid a_{j,k}=0\right]=(\omega_{j}-\sigma_{j})\cdot p^{\prime}_{\text{a},j}-\sigma_{j}\cdot(1-p^{\prime}_{\text{a},j})=\omega_{j}\cdot p^{\prime}_{\text{a},j}-\sigma_{j}=0.$
Similarly, when $a_{j,k}=1$, we have
$\mathbb{E}\left[\tilde{a}_{j,h_{j}(k)}\mid a_{j,k}=1\right]=(\omega_{j}-\sigma_{j})\cdot p^{\prime}_{\text{p},j}-\sigma_{j}\cdot(1-p^{\prime}_{\text{p},j})=\omega_{j}\cdot p^{\prime}_{\text{p},j}-\sigma_{j}=1.$
5.3 Variance of $\tilde{a}_{i,c_{t}}$
@@
We have
$\Pr\left[a^{\prime}_{j,h_{j}(k)}=1\mid a_{j,k}=0\right]=p^{\prime}_{\text{a},j}=\frac{\sigma_{j}}{\omega_{j}}.$
On the other hand,
$\Pr\left[a^{\prime}_{j,h_{j}(k)}=1\mid a_{j,k}=1\right]=p^{\prime}_{\text{p},j}=\frac{1+\sigma_{j}}{\omega_{j}}.$
In Lemma 2 and later, we work under the assumption that $d_{i}<m\approx n/s$. This assumption is reasonable for a large class of graphs as the maximum degree can often be inherently limited by physical or practical constraints. For instance, if we consider a graph representing physical interactions over the past two weeks–relevant in contexts such as epidemic modeling–the number of interactions any individual can have is bounded and cannot approach the size of the global population. Another example is the case of Facebook that imposes a hard cap of 5,000 friends per user Chandler (2012). Additionally, given that the objective of this article is to mitigate the increase in communication overhead caused by privacy measures, this assumption aligns well with our intended use-case. Notably, should $d_{i}$ exceed $m$, the communication cost associated with our method actually becomes lower than that of the corresponding non-private algorithm. To better under the consequence on the accuracy when that assumption is violated, we conducted experiments on power-law graphs in Section 7.1.5.
@@
From the relation $\varepsilon^{\prime}=\ln(1+s(e^{\varepsilon_{1}}-1))$, it follows that
$e^{\varepsilon^{\prime}}-1=s(e^{\varepsilon_{1}}-1).$
Therefore, we obtain
$\omega_{i}=\frac{2+s(e^{\varepsilon_{1}}-1)}{s(e^{\varepsilon_{1}}-1)}\cdot\frac{ms-1}{m-1}=O\left(\frac{s}{e^{\varepsilon_{1}}-1}\right).$
Moreover, under the assumption that $d_{i}<m$, we have
$\sigma_{i}\leq\frac{1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms-1}{m-1}+1=\frac{1}{s(e^{\varepsilon_{1}}-1)}\cdot\frac{ms-1}{m-1}+1\leq 1+\frac{m}{m-1}\cdot\frac{1}{e^{\varepsilon_{1}}-1}=O\left(\frac{1}{e^{\varepsilon_{1}}-1}\right).$
@@
We begin by analyzing the variance of the binary variable $a^{\prime}_{j,h_{j}(k)}$. Since it takes values in $\{0,1\}$, we apply Lemma 4 to obtain
$\text{Var}\left(a^{\prime}_{j,h_{j}(k)}\right)=\Pr\left[a^{\prime}_{j,h_{j}(k)}=1\right]\cdot\left(1-\Pr\left[a^{\prime}_{j,h_{j}(k)}=1\right]\right)=\left(\frac{a_{j,k}+\sigma_{j}}{\omega_{j}}\right)\cdot\left(1-\frac{a_{j,k}+\sigma_{j}}{\omega_{j}}\right)\leq\frac{1+\sigma_{j}}{\omega_{j}}.$
Next, using Lemma 5, we bound the variance of the estimator $\tilde{a}_{j,h_{j}(k)}$ as follows:
$\text{Var}\left(\tilde{a}_{j,h_{j}(k)}\right)=\text{Var}\left(\omega_{j}a^{\prime}_{j,h_{j}(k)}+\tilde{\sigma}_{i}\right)=\omega_{j}^{2}\text{Var}\left(a^{\prime}_{j,h_{j}(k)}\right)+\text{Var}\left(\tilde{\sigma}_{i}\right)\leq\omega_{j}(1+\sigma_{j})+\left(\frac{s-1}{ms-s}\right)^{2}\cdot\frac{2}{\varepsilon_{0}^{2}}.$
Therefore, we conclude, using the assumption that for all $j$, $d_{j}<m$ and Lemma 5, that
$\text{Var}\left(\tilde{a}_{j,h_{j}(k)}\right)=O\left(\frac{s}{(e^{\varepsilon_{1}}-1)^{2}}+\frac{s^{2}}{n^{2}\varepsilon_{0}^{2}}\right).$
It follows from Theorem 5 and subsequent theorems that the variance component associated with $\varepsilon_{1}$ dominates the component associated with $\varepsilon_{0}$. The $\varepsilon_{0}$-dependent term scales as $1/n^{2}$ and is therefore negligible on large networks. Under a fixed privacy budget $\varepsilon=\varepsilon_{0}+\varepsilon_{1}+\varepsilon_{2}$, this suggests setting $\varepsilon_{0}\ll\varepsilon_{1}$ to minimize variance. On the other hand, the sensitivity of the counting step—and the error introduced at Line 5 (Laplace or smooth sensitivity)—can be large for certain networks; accordingly, we recommend choosing $\varepsilon_{1}\approx\varepsilon_{2}$.
In Section 5.5, we will elaborate that $\tilde{a}_{j,h_{j}(k)}$ can cut the communication cost by a factor of $s^{2}$ in contrast to using the direct randomized response results $a^{\prime}_{j,k}$. Given that the variance of $a^{\prime}_{j,k}$ is known to be $\Theta\left(\frac{1}{(e^{\epsilon_{1}}-1)^{2}}\right)$, the previous theorem suggests that, while we manage to decrease the communication cost by a factor of $s^{2}$, this is at the expense of amplifying the variance by a factor of $s$. In the upcoming section, we will demonstrate that, for a set communication cost, this mechanism yields a considerably reduced variance.
5.4 Covariance of Variables
In situations where the goal is to sum up specific estimators, like in subgraph counting, a covariance emerges between these estimators. The size of this covariance is addressed in the subsequent theorem.
@@ 5.4 Covariance of Variables
When $j\neq j^{\prime}$, the two estimators are computed from disjoint sets of random variables and are therefore independent.
@@
When $j=j^{\prime}$, the estimators $\tilde{a}_{j,h_{j}(k)}$ and $\tilde{a}_{j,h_{j}(k^{\prime})}$ are identical if $k$ and $k^{\prime}$ fall into the same bin, i.e., $h_{j}(k)=h_{j}(k^{\prime})$. We denote this event by $C^{(j)}_{k,k^{\prime}}$, which occurs with probability
$\Pr\left(C^{(j)}_{k,k^{\prime}}\right)=\frac{s-1}{ms-1}.$
As a result, for $k,k^{\prime}<j$,
@@
$\operatorname{Cov}\left(\tilde{a}_{j,h_{j}(k)},\tilde{a}_{j,h_{j}(k^{\prime})}\right)=\omega_{j}^{2}\cdot\operatorname{Cov}\left(a^{\prime}_{j,h_{j}(k)},a^{\prime}_{j,h_{j}(k^{\prime})}\right)=\omega_{j}^{2}\left(\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\right]-\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}\right]\mathbb{E}\left[a^{\prime}_{j,h_{j}(k^{\prime})}\right]\right).$
@@
We consider the conditional expectations $\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\mid C^{(j)}_{k,k^{\prime}}\right$] and $\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\mid\overline{C^{(j)}_{k,k^{\prime}}}\right$].
Recall from the previous section that
$\Pr\left[a_{j,h_{j}(k)}=1\right]=\frac{s-1}{s}\cdot\frac{d_{j}}{ms-1}+a_{j,k}\left(\frac{1}{s}-\frac{1}{ms-1}\right).$
Multiplying both sides by $\frac{ms-1}{ms-2}$ yields
$\frac{ms-1}{ms-2}\cdot\Pr\left[a_{j,h_{j}(k)}=1\right]=\frac{s-1}{s}\cdot\frac{d_{j}}{ms-2}+a_{j,k}\left(\frac{ms-1}{s(ms-2)}-\frac{1}{ms-2}\right)\geq\frac{s-1}{s}\cdot\frac{d_{j}-a_{j,k}}{ms-2}+\frac{a_{j,k}}{s}.$
Let $p_{k}=\Pr\left[a_{j,h_{j}(k)}=1\mid\overline{C^{(j)}_{k,k^{\prime}}}\right$]. Then,
$p_{k}=\frac{a_{j,k}}{s}+\frac{s-1}{s}\cdot\frac{d_{j}-a_{j,k}-a_{j,k^{\prime}}}{ms-2}\leq\frac{ms-1}{ms-2}\cdot\Pr\left[a_{j,h_{j}(k)}=1\right].$
It follows that
$\Pr\left[a^{\prime}_{j,h_{j}(k)}=1\text{ and }a^{\prime}_{j,h_{j}(k^{\prime})}=1\mid\overline{C^{(j)}_{k,k^{\prime}}}\right$] $=\left(p_{k}\cdot\frac{e^{\varepsilon^{\prime}}-1}{e^{\varepsilon^{\prime}}+1}+\frac{1-p_{k}}{e^{\varepsilon^{\prime}}+1}\right)\times\left(p_{k^{\prime}}\cdot\frac{e^{\varepsilon^{\prime}}-1}{e^{\varepsilon^{\prime}}+1}+\frac{1-p_{k^{\prime}}}{e^{\varepsilon^{\prime}}+1}\right)$
$\leq\left(\frac{ms-1}{ms-2}\right)^{2}\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}\right]\mathbb{E}\left[a^{\prime}_{j,h_{j}(k^{\prime})}\right].$
Hence, for $s\geq 3$ (which is always satisfied in practice),
$\Pr\left[\overline{C^{(j)}_{k,k^{\prime}}}\right]\cdot\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\mid\overline{C^{(j)}_{k,k^{\prime}}}\right]-\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}\right]\mathbb{E}\left[a^{\prime}_{j,h_{j}(k^{\prime})}\right]\leq 0.$
Thus,
@@
$\operatorname{Cov}\left(\tilde{a}_{j,h_{j}(k)},\tilde{a}_{j,h_{j}(k^{\prime})}\right)\leq\omega_{j}^{2}\cdot\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\mid C^{(j)}_{k,k^{\prime}}\right]\cdot\Pr\left(C^{(j)}_{k,k^{\prime}}\right)$.
Let $p_{k}^{\prime}=\Pr\left[a_{j,h_{j}(k)}=1\mid C^{(j)}_{k,k^{\prime}}\right$]. Then,
$p_{k}^{\prime}=\frac{a_{j,k}+a_{j,k^{\prime}}}{s}+\frac{s-2}{s}\cdot\frac{d_{j}-a_{j,k}-a_{j,k^{\prime}}}{ms-2}\leq\Pr\left[a_{j,h_{j}(k)}=1\right]+\Pr\left[a_{j,h_{j}(k^{\prime})}=1\right].$
It follows that
@@
Using Lemma 4 and $\Pr\left(C^{(j)}_{k,k^{\prime}}\right)=\frac{s-1}{ms-1}$, we obtain
$\operatorname{Cov}\left(\tilde{a}_{j,h_{j}(k)},\tilde{a}_{j,h_{j}(k^{\prime})}\right)\leq 2\cdot\frac{s-1}{ms-1}\cdot\omega_{j}(1+\sigma_{j})=O\left(\frac{s^{2}}{n(e^{\varepsilon_{1}}-1)^{2}}\right).$
When $s\ll n$, the covariance in the previous theorem is significantly smaller than the variance calculated in Theorem 4. While the variance of the summation includes more covariance terms than variance terms, we will demonstrate in the next section that the contribution from the covariance terms is not substantially larger than the that of the variance terms in Theorem 4.
5.5 Communication Cost